5. A start-up company in Dublin offers online training. They advertise their courses via Google. In 2014 they advertised 30 c please show clearly all the calculations and indicate if you have used statisical tables (and which specific table used, eg. “percentage points of the x2 distribution), thank you!
Answer #1

(a) Standard error = s/\sqrt{n}

For 2014 , n=30 , s= 4.6

Standard error = 0.8398

For 2015 , n=32 , s= 6.1

Standard error = 1.0783

Since standard error is obtained by diving standard deviation by square root of sample size , therefore standard error are smaller.

(b) (1- a)100 % confidence interval for difference in mean number of students in 2014 and 2015 means we are (1- a)100 % confident that the actual difference between mean number of students in two years lies in the interval , which is given by

(\bar{x1}-\bar{x2})\pm t_{c}*\sqrt{s1^{2}/n1+s2^{2}/n2}

where tc is the critical value corresponding to (1- a)100 % confidence , For higher confidence tc is higher , resulting in wider confidence interval.

When the sample sizes is larger , margin of error t_{c}*\sqrt{s1^{2}/n1+s2^{2}/n2} becomes smaller , resulting in narrow / small confidence interval.

(c) Calculation of df

(s 12/n 1 + s22/n2)2 = 3.49

(12/n1)2/011-1) + (s22/n2)2/(n2-1) = 0 060S

df = 57

tc = 2.002 (from t critical value table )

therefore 95% confidence interval is

(44.2-41.5) ± 2.002 * V4.62/30 +6.12/32

2.7± 2.74

=(-0.04 , 5.44)

(d) key assumptions are

(i) population are normally distributed

(ii) observations are independent between and within samples

Since it is not mentioned whether populations are normally distributed , this assumption might be violated , which can be fixed by taking large samples .